The closest distance a book can be read from a pair of reading eyeglasses (Power = 1.96 dp) is 26.5 cm. What is the near distance?(Assume a distance between the eyeglasses and the eyes to be 3.00 cm)
Answer #1

given

distance between book to eye is

do = 26.5 + 3

= 29.5 cm

do = 0.295 cm

the equation is using here 1/f = 1/do + 1/di

1/0.295 + 1/di = 1/f

D = 1/f

= 1.96

1.96 = 1/0.295 + 1/di

1/di = 1.96 - (1/0.295)

1/di = - 1.4298

di = - 0.6993 m

or

so the near distance is di = 69.93 cm